OpenStudy (anonymous):

Simplify (x + 2)2 – (x -7)(2x + 3) and show steps.

OpenStudy (angela210793):

(x + 2)2 – (x -7)(2x + 3) x^2+4x+4-(x^2+3x-14x-21)=x^2+4x+4-x^2+11x+21=15x+25